### v1

# Exploring a tentative link between FRBs and pulsars with broad energy distributions and applications for nearby FRB survey strategies

Abstract
Fast radio bursts (FRBs) are energetic, short-duration radio pulses of unclear origins. In this study, we investigate the FRBs and pulsars with broad energy distributions by fitting their high energy tails with a power-law model. Two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB (FRB 20200120E), and two pulsars (RRATs J1846--0257 and J1854+0306), exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their bright pulses contribute significantly to the total radio pulse energy. The brightest bursts from these sources fit well with a simple power-law model ($\alpha=-0.26\pm 0.05$), indicating a tentative link between certain high-luminosity FRBs and low-luminosity radio bursts. We also discuss detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF in the search for FRBs in nearby globular clusters (GCs) using different power-law indices, recommending targets for observation. We suggest that combining observations with FAST ($\sim 3$ hours) and Parkes cryoPAF (10–20 hours) are practicable for discovering new FRBs in the nearby GCs.
Radio bursts (1339), Radio transient sources (2008), Radio pulsars (1353)
1 Introduction
Fast radio bursts (FRBs) are immensely energetic radio pulses with durations ranging from microseconds to seconds (Lorimer et al., 2007; Snelders et al., 2023; Chime/Frb Collaboration et al., 2022). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeat bursts (CHIME/FRB Collaboration et al., 2021; Spitler et al., 2016; Chime/Frb Collaboration et al., 2023). The origin of FRBs remains unclear, but the only two nearby events offer clues to their engines: the apparently one-off burst FRB 20200428 from the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) and the repeating FRB 20200120E, localized to a globular cluster (GC) in the M81 galactic system at a distance of 3.63 Mpc (Bhardwaj et al., 2021; Kirsten et al., 2022). While these two events suggest connections to distinct environments, young magnetars and ancient star populations, respectively, both have propelled FRB theories involving compact objects, especially neutron stars, as potential sources (Zhang, 2023).
The typical energy of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020) is approximately four orders of magnitude greater than that of the two nearby FRBs (Bochenek et al., 2020; Nimmo et al., 2023), and about twelve orders of magnitude larger than that of pulsars [^1] http://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005). However, the detection of a bright burst from FRB 20200120E (Zhang et al., 2024a) bridges the gap between nearby and cosmological FRBs. Similarly, the detection of radio bursts with varying luminosities from SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Zhu et al., 2023) connects the energy range of nearby FRBs to that of pulsars. As presented in Figure 1 of Zhang et al. (2024a), no clear gap exists between the estimated isotropic-equivalent energy releases of various radio pulses.
It is intriguing to explore whether nearby radio pulse phenomena can be connected to cosmological FRBs, particularly repeating ones. Follow-up observation of repeating FRBs (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023) have revealed extremely variable properties, especially in energy, spanning more than three orders of magnitude. In contract, broad energy distributions are atypically for single pulses from pulsars (Burke-Spolaor et al., 2012). Intensity variations exceeding an order of magnitude are usually attributed to the rare phenomenon of “giant pulses” (Kuzmin, 2007), reported in only about ten (Wang et al., 2019) of the more than 3500 known pulsars (Manchester et al., 2005). However, recent observations of some of another kind of special pulsars (Zhang et al., 2024b, 2024), rotating radio transients (RRATs), characterized by sporadic radio pulses, have also shown broad energy distributions. It is still worthy to explore the potential relation between FRBs and these pulsars with broad energy distributions.
A few hyperactive repeating FRBs provide large samples of $\gtrsim 1000$ burst, enabling detailed studies of their energy distributions (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023). Notably, two such sources–FRB 20201124A and FRB 20220912A–have been observed for more than 2000 hours, revealing breaks in their burst energy distributions, with a flattening of the power-law slope ($\gtrsim-1$) at higher energies (Kirsten et al., 2024; Ould-Boukattine et al., 2024). It is also noteworthy that the detection of a bright burst from the nearby FRB 20200120E indicates a similar flat energy distribution, with a power-law index of approximately -0.98 (Zhang et al., 2024a). Furthermore, two RRATs have also exhibited power-law-like tails at high energies with similarly flat indices (Zhang et al., 2024). These observations suggest that flat power-law tails at high energies may offer a clue into the mechanisms underlying these short-duration radio pulses with diverse energy releases.
In this study, we present an analysis of the energy distributions tails at high energy of FRBs and pulsars using a consistent method. Section 2 describes the sources selection and data processing, while Section 3 and 4 presents and discusses the results. We conclude in Section 5.
2 Sources selection and data processing
2.1 Sources selection
Figure 1: Isotropic-equivalent spectral luminosity versus the product of observing frequency and pulse width for various pulsars and FRBs. The brightest strong pulses from RRATs with relatively high $B_{\rm s}$ (J1846-0257 and 1854+0306) are represented by red and black filled stars, while their weak emissions are denoted by unfilled stars (Zhang et al., 2024). The brightest giant pulses from J0534+2200 (Crab pulsar) (Bera & Chengalur, 2019) and J0540--6919  (Geyer et al., 2021) are shown as blue and violet filled triangles, with their normal emissions  (Manchester et al., 2005) marked by unfilled triangles. The brightest burst from the nearby FRB 20200120E (Zhang et al., 2024a) is indicated by an orange filled square, with its relatively weaker bursts (Nimmo et al., 2023) are marked by unfilled squares. The brightest bursts from cosmological FRB 20201124A (Kirsten et al., 2024) and FRB 20220912A (Ould-Boukattine et al., 2024) are represented by black and gray filled circles, with their relatively weaker bursts (Xu et al., 2022; Zhang et al., 2023) shown as unfilled circles. The pulsar (Manchester et al., 2005) population is indicated in thistle. Grey dashed lines represent constant isotropic-equivalent energy release, while grey dotted lines indicate constant brightness temperature.
The selected sources, presented in Figure 1, show the brightest strong pulses that are at least 2--3 orders of magnitude more energetic than their relatively weak emissions. The energies of the detected radio pulses from these sources span roughly 15 orders of magnitude. Specifically, the studied sources include: (1) FRB 20201124A and FRB 20220912A: These are the only two FRBs with up to 2000 hours of follow-up observations, during which ultra-luminous bursts with isotropic energy up to 1041 erg were detected (Kirsten et al., 2024; Ould-Boukattine et al., 2024), placing them among the most energetic FRBs observed. (2) FRB 20200120E: A nearby FRB that exhibited a burst storm (Nimmo et al., 2023) and an abnormal burst more than 42 times brighter than previously detected bursts at a similar frequency (Zhang et al., 2024a). (3) J0534+2200 (the Crab pulsar) and J0540--6919: These pulsars are known for producing giant pulses with the largest excesses relative to their normal emissions (Bera & Chengalur, 2019; Geyer et al., 2021; Wang et al., 2019). (4) RRATs J1846--0257 and J1854+0306: These RRATs have relatively high surface magnetic fields ($\sim 2\times 10^{13}$ G) and show weak emissions that are up to 687 and 512 times fainter than their strong individual pulses. We also analysed data from RRAT J1913+1330 (Zhang et al., 2024b), which exhibits extreme pulse-to-pulse modulation and an energy distribution spanning 3 orders of magnitude but could be well-fitted by two log-normal functions. Its distribution tail at the high energy part provides a useful comparison with the four source kinds above.
2.2 Data processing
To compare the energy distributions of different radio pulses, we calculated the specific isotropic-equivalent energy using (Zhang, 2018):
$\centering E_{S}=\frac{4\pi D_{L}^{2}}{1+z}\,F,\@add@centering$ (1)
where $D_{L}$ is the luminosity distance, $z$ is the redshift, and $F$ is the pulse fluence. The energy values for each source were then normalized to their brightest bursts.
Since only one bright burst has been detected from FRB 20200120E (Zhang et al., 2024a) and detailed burst properties for FRB 20220912A are currently unavailable (Ould-Boukattine et al., 2024), we were unable to reprocess their high-energy distributions. To compare our results with their reported distributions, which are cumulative, our initial analysis used a cumulative distribution, adjusting the bins for each source to cover a similar range in log-10 space. We also provide a differential distribution analysis. To enable comparisons between different sources, we defined a density that is calculated as the number of bursts per bin divided by the total number of detected bursts.
The power-law-like tails were fitted using the Python package curve_fit function from the SciPy library (Virtanen et al., 2020). For data with complex uncertainties, we performed the fitting within a Bayesian framework using Markov Chain Monte Carlo (MCMC) sampling, via the emcee Python package (Foreman-Mackey et al., 2013).
3 Results
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Cumulative specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. Each high-energy tail is fitted with a simple power law model.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Differential specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. Each high-energy tail is fitted with a simple power law model.
Figure 4: Comparison of cumulative (top) and differential (bottom) normalized specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919. The energy values for each source are normalized to their respective brightest bursts. The density represents the number of bursts per bin divided by the total number of detected bursts.
Figure 2 presents the cumulative specific energy distributions of high-energy bursts from the selected sources with relatively sufficient burst counts, most of which can be well-fitted by a simple power-law model. The high-energy tail of the distribution for RRAT J1913+1330 is forcefully fitted using a power-law function, although the fit is not as good as for the other sources, in order to enable comparison. Specifically, the fitted power-law indices ($\alpha$) for FRB 20201124A (46 bursts  [^2] For events detected simultaneously by multiple telescopes, we selected bursts from the observation with the largest bandwidth.), RRATs J1846--0257 (72 bright pulses) and J1854+0306 (325 bright pulses) are $-0.75\pm 0.14$, $-0.84\pm 0.10$ and $-0.76\pm 0.03$, respectively, all of which are $\gtrsim-1$. In contrast, the indices for the Crab pulsar ($-1.98\pm 0.07$, 1955 giant pulses), J0540--6919 ($-2.29\pm 0.07$, 865 bright single pulses) and RRAT J1913+1330 ($-3.04\pm 0.29$, 1955 individual pulses) are significantly steeper, at $\lesssim-2$. Although the available data for bursts from FRB 20200120E and FRB 20220912A are insufficient to reprocess their high-energy distributions, the reported fitted power-law indices are $-0.98\pm 0.06$ (Zhang et al., 2024a) for FRB 20200120E and range from $-0.99$ to $-0.60$ (Ould-Boukattine et al., 2024) for FRB 20220912A at around 1.4 GHz, both indicate relatively flat indices.
Furthermore, we also present the differential specific energy distributions of high-energy bursts from these source in Figure 3. Similar to the cumulative distributions, the fitted indices using a power-law function for the differential distributions at the high energy tails ($\beta$) of FRB 20201124A ($-0.40\pm 0.21$), RRATs J1846--0257 ($-0.63\pm 0.16$) and J1854+0306 ($-0.65\pm 0.05$) are significantly flatter than those of the Crab pulsar ($-1.98\pm 0.09$), J0540--6919 ($-2.0\pm 0.1$) and RRAT J1913+1330 ($-2.72\pm 0.31$). However, the sample sizes for some sources, especially RRAT J1846--0257 and FRB 20201124A, are limited, and the fitted uncertainties are relatively large.
As shown in the comparison in Figure 4, the energy distributions of FRB 20201124A, RRATs J1846--0257 and J1854+0306 decline much more slowly at high energies compared to those of the giant-pulse pulsars and RRAT J1913+1330.
4 Discussion
4.1 Bright bursts of various radio sources
Figure 5: Specific energy of the brightest bursts versus their detection rates for FRB 20201124A, FRB 20220912A, FRB 20200120E and the RRATs J1846--0257 and J1854+0306. The data are fitted with a power law relation (red dashed line) using a Bayesian framework, and the shaded region represents the 68% credible interval for the fit. Posterior distributions and a parameter correlation plot for the power law fit parameters are shown in the inset.
Sources with flat energy distributions ($\alpha\gtrsim-1$) at high energies suggest that their “bright pulses” might contribute significantly to the total radio pulse energy release. Therefore, the detection of the brightest bursts from these sources could provide insight into a potential connection between cosmological high-luminosity radio bursts and nearby low-luminosity ones. Figure 5 presents the specific energy of the brightest bursts and the corresponding detection rates for the studied sources with flat energy distributions, i.e., RRATs J1846--0257 and J1854+0306, FRB 20200120E, FRB 20201124A and FRB 20220912A. Interestingly, these data points align well with a simple power-law fit of $\alpha=-0.26\pm 0.05$. Due to the sample size of only five points and the stochastic nature of detecting the “brightest burst”, we remain cautious about interpreting this correlation as a physical relationship. Nevertheless, this trend suggests that high-luminosity radio bursts, such as FRBs, could be naturally derived from low-luminosity radio bursts, such as those from pulsars. Although an extremely shallow power-law index is implied, the flatter energy distributions at higher energies observed in these sources may provide clues for understanding this tentative connection.
4.2 Survey strategies for nearby FRBs
As presented in Figure 2 and 3 of Section 3, the fitting of the cumulative energy distribution at high energies using a power-law function yields better results with smaller uncertainties compared to the differential distributions. This difference may arise because the limited burst numbers for the studied sources are insufficient to provide a reliable sample for the differential distributions. Therefore, we used the power-law indices ($\alpha$) derived from the cumulative energy distribution to estimate the detection rate in this section, which is given by:
$\centering{\rm log}_{10}N(\geq E)=C+\alpha\,{\rm log}_{10}E,\@add@centering$ (2)
where $N(\geq E)$ represents the predicted number of bursts with specific energy above $E$.
For a telescope operating at a fluence completeness threshold of $F_{\rm det}$, bursts with energy above the threshold $E_{\rm det}=4\pi D_{\rm det}^{2}F_{\rm det}$ are expected to be detected, where $D_{\rm det}$ is the distance to the target source, and assuming a detected burst rate of $R_{\rm det}$. Applying Equation 2, the burst rate expected for bursts detectable by an arbitrary radio telescope with a fluence threshold $F_{\rm th}$, for a source at a distance of $D_{\rm tar}$, can be written as:
$\centering R_{\rm th}\approx R_{\rm det}\times\left(\frac{F_{\rm th}}{F_{\rm det}}\right)^{\alpha}\left(\frac{D_{\rm tar}}{D_{\rm det}}\right)^{2\alpha}.\@add@centering$ (3)
Kremer et al. (2023) presented prospects for detecting FRBs from GC in nearby galaxies using FAST (Jiang et al., 2020) and MeerKAT (Bailes et al., 2020), suggesting that NGC 4486 (M87) could be the best candidate target. However, the fluence threshold ($F_{\rm th}$) of a telescope may be significantly affected by the continuum radio flux densities ($S_{\rm tar}$) from the target galaxies (Suresh et al., 2021), which was not thoroughly discussed in detection rate estimation by Kremer et al. (2023). Taking into account $S_{\rm tar}$ and the telescope’s corresponding system-equivalent flux density (SEFD), the radiometer equation (Lorimer & Kramer, 2004) for estimating the minimum detectable fluence can be written as:
$F_{\rm min}=\frac{{\rm S/N}_{\rm min}({\rm SEFD}+S_{\rm tar})}{\sqrt{{\Delta}{\nu}N_{p}W}}\times W,$ (4)
where we ignore the loss factor, ${\rm S/N}_{\rm min}\sim 7$ is the signal-to-noise ratio (S/N) threshold typically applied in the single-pulse search pipeline, ${\Delta}{\nu}$ is the bandwidth, $N_{p}$ is the number of polarization, and $W$ is the burst width  [^3] In this study, we used $W\sim 1$ ms to estimate the sensitivity.. The FAST telescope has a system temperature of $\sim$ 20 K and a gain of $16{\rm K/Jy}$ (Jiang et al., 2020), resulting in a corresponding SEFD of $S_{\rm FAST}=1.25$ Jy, while MeerKAT has a system temperature of $\sim$ 18 K and a gain of $2.8{\rm K/Jy}$ (Bailes et al., 2020), resulting in a corresponding SEFD of $S_{\rm MeerKAT}=6.43$ Jy. The galaxy M87 could contribute $S_{\rm tar}$ up to 210 Jy at 1.4 GHz (Brown et al., 2011), meaning the effective detectable fluence of sources in M87 for FAST and MeerKAT would be approximately 169 and 34 times larger, respectively. Therefore, for telescopes with high sensitivities, galaxies with relatively small $S_{\rm tar}$, such as NGC 4649 (M60), which has $S_{\rm tar}\sim 0.03$ at 1.4 GHz (Brown et al., 2011) and is estimated to contain approximately 4000 GCs (Harris et al., 2013), could be more efficient search targets.
At a given distance, the number of potential sources is proportional to the field of view (FOV) of the telescope. Combining the FOV with Equation 3, the detection rate of the telescope is given by:
$\centering R\propto F_{\rm th}^{\alpha}\cdot D_{\rm tar}^{2\alpha}\cdot{\rm FOV}.\@add@centering$ (5)
For smaller values of $\alpha$, the telescope’s sensitivity dominates the detection rate, whereas for larger values of $\alpha$ the FOV becomes more important. Therefore, in addition to telescopes with high sensitivity, such as FAST and MeerKAT, observations with large FOV but relatively lower sensitivity may also be efficient for capturing such events.
Figure 6: Parkes cryoPAF 63 pointings covering more than 10,000 valid GCs. The two most suggested pointings, TP0 (Dec $\sim$ 13 deg) and TP1 (Dec $\sim$ 8 deg), are marked by black and red stars, respectively, with their corresponding sky coverage shown by circles.
Table 1: Properties of the 24 GC systems covered by Parkes cryoPAF pointing TP0 (RA = 187.2 deg; Dec = 13.2 deg). $d_{\rm L}$ and $D_{\rm TP}$ represent the distance from the GC system centre to the earth and the pointing, respectively. References for the galaxy continuum radio flux densities at 1.4 GHz ($S_{\rm tar,1.4GHz}$) are: (a) Brown et al. (2011), (b) Allison et al. (2014), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/), (d) Nyland et al. (2017), (e) Oosterloo et al. (2010).
Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm V\_GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
NGC4486 (M87) 187.706 12.391 17.0 0.95 13000 6500 $\sim$ 210 (a)
NGC4374 (M84) 186.265 12.887 18.5 0.96 4301 4301 $\sim$ 6.1 (b)
NGC4406 (M86) 186.549 12.946 17.1 0.68 2800 2800 $<$ 0.1 (c)
NGC4473 187.454 13.429 15.2 0.34 376 376 $<$ 0.1 (a)
NGC4435 186.919 13.079 16.6 0.30 345 345 $<$ 0.1 (c)
NGC4459 187.250 13.979 16.0 0.78 218 218 $<$ 0.1 (d)
NGC4474 187.473 14.069 15.5 0.91 116 116 $<$ 0.1 (c)
NGC4458 187.240 13.242 16.3 0.06 72 72 $<$ 0.1 (e)
NGC4387 186.424 12.810 18.0 0.85 70 70 $<$ 0.1 (e)
VCC-940 186.696 12.454 18.7 0.89 61 61 $<$ 0.1 (c)
NGC4479 187.577 13.578 17.4 0.53 59 59 $<$ 0.1 (c)
NGC4478 187.574 12.327 17.0 0.95 58 58 $<$ 0.1 (c)
VCC-1386 187.964 12.657 16.0 0.92 27 27 $<$ 0.1 (c)
NGC4476 187.496 12.349 17.7 0.90 20 20 $<$ 0.1 (c)
VCC-871 186.523 12.560 16.0 0.92 18 18 $<$ 0.1 (c)
VCC-965 186.763 12.561 16.9 0.77 15 15 $<$ 0.1 (c)
VCC-1185 187.348 12.451 16.9 0.76 14 14 $<$ 0.1 (c)
NGC4431 186.864 12.290 15.8 0.97 11 11 $<$ 0.1 (c)
VCC-1104 187.117 12.824 15.7 0.38 8 8 $<$ 0.1 (c)
VCC-896 186.594 12.787 16.0 0.72 5 5 $<$ 0.1 (c)
VCC-1105 187.114 14.156 16.0 0.96 5 5 $<$ 0.1 (c)
NGC4486B 187.633 12.490 16.3 0.83 4 4 $\sim$ 0.4 (c)
VCC-1272 187.567 13.308 16.0 0.37 3 3 $<$ 0.1 (c)
VCC-1077 187.043 12.807 16.0 0.42 2 2 $<$ 0.1 (c)
Table 2: Properties of the 12 GC systems covered by Parkes cryoPAF pointing TP1 (RA = 186.6 deg; Dec = 8.1 deg). $d_{\rm L}$ and $D_{\rm TP}$ represent the distance from the GC system centre to the earth and the pointing, respectively. References for the galaxy continuum radio flux densities at 1.4 GHz ($S_{\rm tar,1.4GHz}$) are: (a) White & Becker (1992), (b) Brown et al. (2011), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/).
Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm V\_GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
NGC4472 (M49) 187.445 8.000 17.0 0.84 7000 7000 $\sim$ 0.8 (a)
NGC4365 186.118 7.318 23.3 0.92 3246 3246 $<$ 0.1 (b)
NGC4434 186.903 8.154 22.5 0.30 141 141 $<$ 0.1 (c)
VCC-1254 187.521 8.074 16.0 0.91 26 26 $<$ 0.1 (c)
VCC-1107 187.126 7.325 16.0 0.93 25 25 $<$ 0.1 (c)
NGC4464 187.339 8.157 15.8 0.73 25 25 $<$ 0.1 (c)
NGC4318 185.680 8.198 22.1 0.92 18 18 $<$ 0.1 (c)
VCC-747 186.199 8.992 18.5 0.98 16 16 $<$ 0.1 (c)
VCC-1167 187.301 7.879 16.0 0.73 14 14 $<$ 0.1 (c)
VCC-571 185.671 7.950 23.8 0.93 11 11 $<$ 0.1 (c)
VCC-1049 186.978 8.090 16.0 0.37 8 8 $<$ 0.1 (c)
VCC-992 186.828 8.213 16.0 0.25 4 4 $<$ 0.1 (c)
A cryogenic phased array receiver (cryoPAF) is currently being completed and tested on the Parkes telescope. The new cryoPAF is expected to have a system temperature of $<$ 20 K and a gain of $0.735{\rm K/Jy}$, resulting in a corresponding SEFD of $S_{\rm Parkes}=27.21$ Jy. Its wide FOV spans up to 3 square degrees and consists of 72 adjacent beams, covering a frequency range of 700-1950 MHz  [^4] As cryoPAF is still being tested, the accurate parameters of the full-beam system performance are unavailable for now. We used the values referenced from the Parkes cryoPAF introduction by Alex Dunning: https://research.csiro.au/ratechnologies/wp-content/uploads/sites/295/2022/11/PAFAR2022-Dunning-CryoPAF_for_Parkes.pdf. Based on the cryoPAF’s sky coverage, which is modelled as a circle with a radius of approximately 1 deg in this work, and the catalogue of GC systems (Harris et al., 2013) within a relatively close distance ($<$ 40 Mpc), we conducted a grid analysis of the sky with a 0.1 deg step size to estimate the number of GCs covered by the cryoPAF FOV. Notably, about 20% of GCs lie within a 4-arcminute radius of the centre of M87 (Strader et al., 2011), where the radio background is comparable to $S_{\rm Parkes}$ (Perley & Butler, 2017). Given this, the cryoPAF observation should cover the majority of GCs in M87 with adequate sensitivity, and we adopt a conservative estimate that 50% of the GCs are valid targets for Parkes observation (denoted as $N_{\rm V\_GC}$). As shown in Figure 6, among the more than 6,000,000 analysed pointings, 63 could cover more than 10,000 valid GCs, all with Dec near 13 or 8 deg. Pointings TP0 (RA = 187.2 deg; Dec = 13.2 deg) and TP1 (RA = 186.6 deg; Dec = 8.1 deg) cover the largest number of valid GCs in these two Dec ranges. Based on this, we suggest these two pointings as the best targets for searching for FRBs in nearby GCs. As listed in Tables 1 and 2, the Parkes cryoPAF pointing TP0 could cover 24 GC systems, expected to include 21,607 GCs, of which 15,107 are valid targets. TP1 could cover 12 GC systems, expected to include 10,535 GCs, all of which are valid.
Table 3: Predicted FRB detection rates in nearby GC systems by FAST, MeerKAT and Parkes cryoPAF.
FRB detection rate (hr-1)
Target Instrument $N_{\rm GC}$ $N_{\rm V\_GC}$ $\alpha$ = --0.2 $\alpha$ = --0.5 $\alpha$ = --1.0 $\alpha$ = --1.5 $\alpha$ = --2.0
M60 FAST 4000 1203 0.16(0.01--0.48) 0.35(0.02--1.06) 1.36(0.07--4.08) 5.21(0.27--15.64) 19.98(1.02--59.93)
M60 MeerKAT 4000 4000 0.37(0.02--1.11) 0.49(0.02--1.46) 0.77(0.04--2.31) 1.21(0.06--3.64) 1.92(0.1--5.75)
TP0 Parkes cryoPAF 21607 15107 1.3(0.07--3.89) 1.53(0.08--4.6) 2.03(0.1 --6.1) 2.72(0.14--8.15) 3.65(0.19--10.96)
TP1 Parkes cryoPAF 10535 10535 0.88(0.04--2.65) 1.01(0.05--3.03) 1.28(0.07--3.84) 1.65(0.08--4.94) 2.15(0.11--6.44)
[…]
As presented in Table 3, for steep power-law indices ($\alpha=-1.5,\,-2.0$), FAST is the most powerful telescope for detecting nearby FRBs, with a detection rate of up to 20 hr-1. Even at the lower detection rate limits, only about 3 hours of observation are required to expect a new detection. In contrast, Parkes cryoPAF is more efficient for flatter indices ($\alpha=-0.2,\,-0.5,\,-1.5$), with 10 to 20 hours of observation typically required to ensure detecting a new source.
5 Conclusion
Before the discovery of a population of FRBs (Thornton et al., 2013) that supported the astrophysical origin of the “Lorimer burst” (Lorimer et al., 2007), pulsars (Manchester et al., 2005) were considered the primary sources of bright, short-duration ($\lesssim$ second) radio pulses, which require coherent emission (Pietka et al., 2015). While FRBs also necessitate coherent emission, their luminosities are more than ten orders of magnitude higher. Initially, the apparent one-off nature of FRBs led to favouring catastrophic scenarios (Zhang, 2023), distinguishing them from pulsars. However, the discovery of repeating FRBs (Spitler et al., 2016), and the investigation that some Galactic radio pulsing sources might be one-off (Keane, 2016; Zhang & Yang, 2024), have reduced this distinction. Further observations of fainter FRB bursts (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Nimmo et al., 2023; Zhang et al., 2024a) and brighter pulses from magnetars or pulsars (Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Bera & Chengalur, 2019) suggest no clear gap between the isotropic-equivalent energy releases of various radio pulses (Zhang et al., 2024a). Additionally, polarimetric similarities between FRBs and pulsars have been observed (Feng et al., 2022; Niu et al., 2024). These findings highlight the importance of exploring potential connections between FRBs and pulsars, at least for certain sources.
In this study, we have presented similarities in the high-energy distributions of two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB 20200120E, and two pulsars (RRATs J1846--0257 and J1854+0306). All these sources exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their “bright pulses” might significantly contribute to the total radio pulse energy release. Furthermore, the brightest bursts from these sources fit well with a simple power-law model of $\alpha=-0.26\pm 0.05$. However, considering the limited sample size, we advise caution in interpreting this correlation as physical.
We encourage further observations to expand the sample of high-energy bursts from both FRB-like and pulsar-like sources. Current FRB surveys may be limited by selection effects, potentially missing FRBs with relative low-luminosity from the nearby GC systems (Zhang et al., 2024a). Survey strategies will need to vary for sources with different energy distributions. In this work, we have discussed detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF for FRB in nearby GCs and suggested optimal survey targets. In principle, conducting combining observations for these targets using FAST for about $3$ hours and Parkes cryoPAF for $10-20$ hours could potentially lead to the discovery of new FRBs in nearby GCs, in practice. Additionally, a lack of detection after even a few hours of FAST observation can still provide strict constraints on flat $\alpha$ values.
Acknowledgments
This research has been partially funded by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos. 12041306,12273113,12233002,12003028,12321003), the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-063), the International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the ACAMAR Postdoctoral Fellow, China Postdoctoral Science Foundation (grant No. 2020M681758), and the Natural Science Foundation of Jiangsu Province (grant Nos. BK20210998).

### v2

@@
-Exploring a tentative link between FRBs and pulsars with broad energy distributions and applications for nearby FRB survey strategies
+Flat tails in FRB and pulsar energy distributions: implications for optimizing nearby FRB surveys
 Abstract
-Fast radio bursts (FRBs) are energetic, short-duration radio pulses of unclear origins.
-In this study, we investigate the FRBs and pulsars with broad energy distributions by fitting their high energy tails with a power-law model.
-Two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB (FRB 20200120E), and two pulsars (RRATs J1846--0257 and J1854+0306), exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their bright pulses contribute significantly to the total radio pulse energy.
-The brightest bursts from these sources fit well with a simple power-law model ($\alpha=-0.26\pm 0.05$), indicating a tentative link between certain high-luminosity FRBs and low-luminosity radio bursts.
-We also discuss detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF in the search for FRBs in nearby globular clusters (GCs) using different power-law indices, recommending targets for observation.
-We suggest that combining observations with FAST ($\sim 3$ hours) and Parkes cryoPAF (10–20 hours) are practicable for discovering new FRBs in the nearby GCs.
+Fast radio bursts (FRBs) are energetic, short-duration radio pulses of unclear origin.
+To explore effective survey strategies for detecting FRBs from nearby globular clusters (GCs), we investigate the burst energy distribution, which has a strong influence on the detection rate.
+We re-analyze FRBs and pulsars exhibiting broad energy distributions by fitting their high-energy tails with power-law models.
+Two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB (FRB 20200120E), and two pulsars (RRATs J1846-0257 and J1854+0306), exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their bright pulses contribute significantly to the total radio pulse energy.
+The brightest bursts from these sources can be fitted with a power-law model ($\alpha_{\rm Bri}=-0.26\pm 0.05$), suggesting that an extremely flat index is required to naturally derive high-luminosity FRBs from low-luminosity sources.
+We present detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF in the search for FRBs in nearby GCs using different power-law indices, recommending targets for observation.
+We suggest that combining observations with FAST ($\sim 3$ hours) and Parkes cryoPAF (10-20 hours) is practicable for discovering new FRBs in nearby GCs.
 Radio bursts (1339), Radio transient sources (2008), Radio pulsars (1353)
@@
 Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeat bursts (CHIME/FRB Collaboration et al., 2021; Spitler et al., 2016; Chime/Frb Collaboration et al., 2023).
-The origin of FRBs remains unclear, but the only two nearby events offer clues to their engines: the apparently one-off burst FRB 20200428 from the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) and the repeating FRB 20200120E, localized to a globular cluster (GC) in the M81 galactic system at a distance of 3.63 Mpc (Bhardwaj et al., 2021; Kirsten et al., 2022).
+The origin of FRBs remains unclear, but the only two nearby events offer clues to their engines: the apparently one-off burst FRB 20200428 from the Galactic magnetar SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020) and the repeating FRB 20200120E, localised to a globular cluster (GC) in the M81 galactic system at a distance of 3.63 Mpc (Bhardwaj et al., 2021; Kirsten et al., 2022).
 While these two events suggest connections to distinct environments, young magnetars and ancient star populations, respectively, both have propelled FRB theories involving compact objects, especially neutron stars, as potential sources (Zhang, 2023).
-The typical energy of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020) is approximately four orders of magnitude greater than that of the two nearby FRBs (Bochenek et al., 2020; Nimmo et al., 2023), and about twelve orders of magnitude larger than that of pulsars [^1] http://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005).
-However, the detection of a bright burst from FRB 20200120E (Zhang et al., 2024a) bridges the gap between nearby and cosmological FRBs.
-Similarly, the detection of radio bursts with varying luminosities from SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Zhu et al., 2023) connects the energy range of nearby FRBs to that of pulsars.
-As presented in Figure 1 of Zhang et al.
-(2024a), no clear gap exists between the estimated isotropic-equivalent energy releases of various radio pulses.
-It is intriguing to explore whether nearby radio pulse phenomena can be connected to cosmological FRBs, particularly repeating ones.
-Follow-up observation of repeating FRBs (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023) have revealed extremely variable properties, especially in energy, spanning more than three orders of magnitude.
-In contract, broad energy distributions are atypically for single pulses from pulsars (Burke-Spolaor et al., 2012).
-Intensity variations exceeding an order of magnitude are usually attributed to the rare phenomenon of “giant pulses” (Kuzmin, 2007), reported in only about ten (Wang et al., 2019) of the more than 3500 known pulsars (Manchester et al., 2005).
-However, recent observations of some of another kind of special pulsars (Zhang et al., 2024b, 2024), rotating radio transients (RRATs), characterized by sporadic radio pulses, have also shown broad energy distributions.
-It is still worthy to explore the potential relation between FRBs and these pulsars with broad energy distributions.
-A few hyperactive repeating FRBs provide large samples of $\gtrsim 1000$ burst, enabling detailed studies of their energy distributions (Li et al., 2021; Xu et al., 2022; Zhang et al., 2023).
-Notably, two such sources–FRB 20201124A and FRB 20220912A–have been observed for more than 2000 hours, revealing breaks in their burst energy distributions, with a flattening of the power-law slope ($\gtrsim-1$) at higher energies (Kirsten et al., 2024; Ould-Boukattine et al., 2024).
-It is also noteworthy that the detection of a bright burst from the nearby FRB 20200120E indicates a similar flat energy distribution, with a power-law index of approximately -0.98 (Zhang et al., 2024a).
-Furthermore, two RRATs have also exhibited power-law-like tails at high energies with similarly flat indices (Zhang et al., 2024).
-These observations suggest that flat power-law tails at high energies may offer a clue into the mechanisms underlying these short-duration radio pulses with diverse energy releases.
-In this study, we present an analysis of the energy distributions tails at high energy of FRBs and pulsars using a consistent method.
-Section 2 describes the sources selection and data processing, while Section 3 and 4 presents and discusses the results.
+Nearby FRBs are crucial in understanding FRBs, as they enable precise localisation to identify their surroundings and deep multiwavelength follow-up to resolve their mechanisms.
+Although CHIME, the current most efficient FRB hunter, has been operating since 2018 and can discover FRBs daily (CHIME/FRB Collaboration et al., 2021), only one repeating FRB with a distance $\leq$ 100 Mpc has been detected (Bhardwaj et al., 2021), indicating that a blind survey is challenging for providing a sufficient sample of nearby repeating FRBs.
+As the only nearby repeating FRB 20200120E has been localised to a GC (Kirsten et al., 2022), Kremer et al.
+(2023) presented a dedicated FRB survey of GCs in nearby galaxies, with M87 being identified as the most promising target, where high-sensitivity radio telescopes like FAST or MeerKAT have a 90% probability of detecting a globular cluster FRB within approximately 10 hours of observation.
+However, it is notable that the detection rate of nearby FRBs from such dedicated surveys is significantly influenced by the burst energy distribution (Kremer et al., 2023).
+The nearby radio bursts expected to be detected by high-sensitivity telescopes such as FAST (Jiang et al., 2020) would have relatively low energies, even comparable to pulsar single pulses.
+The expected event rate from surveys such as those proposed by Kirsten et al.
+(2022) assumes a consistent distribution from bright to faint target sources.
+Therefore, it is important to investigate the relationship between bright sources, such as cosmological FRBs, and relatively faint sources, such as nearby FRBs or even bright pulsar single pulses.
+FRB 20200120E stands as the sole nearby repeating FRB at a distance of approximately 3.63 Mpc that enables detailed energy distribution analysis.
+Observations with the Effelsberg telescope recorded 65 bursts (Nimmo et al., 2023; Kirsten et al., 2022) with fluences ranging from 0.04–0.71 Jy ms at 1.4 GHz, including a remarkable “burst storm” event of 53 bursts detected within 40 minutes.
+The burst storm exhibited a steep energy distribution characterized by a power-law tail with index of $\sim-2.39$ (Nimmo et al., 2023).
+However, a recent observation have detected an exceptionally bright burst (fluence $\approx 30$ ms at 1.1-1.7 GHz), 42 times brighter than previous detection at 1.4 GHz, suggesting a flatter high-energy tail with index of approximately 0.98-0.98 (Zhang et al., 2024b).
+The typical energy of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020) is approximately four orders of magnitude greater than that of FRB 20200120E (Nimmo et al., 2023).
+The study of energy distributions in cosmological FRBs began with FRB 20121102A, the first known repeater (Spitler et al., 2016).
+As summarized by Wang & Zhang (2019), the early repeating samples of FRB 20121102A showed cumulative energy distributions of power-law tails with indices from 0.6~-0.6 to 0.8-0.8111It is noteworthy that their definition of the indices differs from that in our study, we have converted their values to match our definition..
+Later analysis with larger samples also indicate flatter high energy tails with indices larger than 1.0-1.0 (Cruces et al., 2021; Hewitt et al., 2022; Jahns et al., 2023; Li et al., 2021).
+Although monitoring of other active repeaters such as FRB 20180916B and FRB 20190520B showed steep indices $($$\alpha<-2.0$) Chime/Frb Collaboration et al., 2020; Zhang et al., 2022), recent studies of two hyperactive repeaters, FRB 20201124A and FRB 20220912A, both detected extreme bright pulses by ultra-long monitoring (>2000>2000 hours), revealing breaks in their burst energy distributions, with a flattening of the power-law slope ($\gtrsim-1$) at higher energies (Kirsten et al., 2024; Ould-Boukattine et al., 2024).
+These flatter indices at high energies mirror FRB 20200120E’s bright burst properties.
+Furthermore, the latest rate analysis of the new detected hyperactive repeater FRB 20240114A also indicates a potential flattening at higher energies (Huang et al., 2025).
+The typical energy of pulsars (Manchester et al., 2005) is about twelve orders of magnitude smaller than that of cosmological repeating FRBs (Li et al., 2021; Xu et al., 2022; Chime/Frb Collaboration et al., 2020).
+The broad energy distributions observed in FRBs are also atypically for single pulses from pulsars (Burke-Spolaor et al., 2012).
+Intensity variations exceeding an order of magnitude are usually attributed to the rare phenomenon of “giant pulses” (Kuzmin, 2007), reported in only about ten (Wang et al., 2019) of the up to 4000 known pulsars [^2] ATNF Pulsar Catalogue v2.6.1: https://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005).
+Pulsar giant pulses normally exhibit steep energy distributions.
+For instance, J0534+2200 (the Crab pulsar) and J0540-6919 are known for producing giant pulses with the largest excesses relative to their normal emissions (Bera & Chengalur, 2019; Geyer et al., 2021; Wang et al., 2019).
+The Crab pulsar (J0534+2200) demonstrates a single power-law distribution with $\alpha\sim-3.0$ across three decades of fluence.
+Similarly, J0540-6919 shows a cumulative distribution with $\alpha=-2.75\pm 0.02$.
+However, recent observations of another kind of special pulsars (Zhang et al., 2024c, a), rotating radio transients (RRATs), characterized by sporadic radio pulses, have also shown broad energy distributions.
+Although one of which (J1913+1330) present steep tails (Zhang et al., 2024c), two other RRATs (J1846-0257 and J1854+0306) have also exhibited power-law-like tails at high energies with similarly flat indices ($\gtrsim-1$; Zhang et al., 2024a).
+These indices match the flatter high-energy regimes of FRB 20200120E and cosmological repeaters like FRB 20201124A and FRB 20220912A.
+The detection of a bright burst from FRB 20200120E (Zhang et al., 2024b) bridges the gap between nearby and cosmological FRBs, and the detection of radio bursts with varying luminosities from SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Zhu et al., 2023) connects the energy range of nearby FRBs to that of pulsars.
+No clear gap exists between the estimated isotropic-equivalent energy releases of various radio pulses (Zhang et al., 2024b), making it natural to explore the burst energy distribution relationship across radio sources with various energy releases.
+Recent studies have shown that flat tails in pulse energy distributions are observed across cosmological FRBs, the nearby FRB 20200120E, and certain pulsars.
+However, analysis of different sources has employed various methods: some studies derived power-law indices from cumulative distributions, while others used differential distributions.
+Energy calculations have varied depending on whether they incorporated central frequency, bandwidth, or burst frequency extent scaling.
+Additionally, the choice of distribution bins and fitting tools can affect the obtained index values.
+In this study, we present a consistent analysis of the high-energy tails in the energy distributions of FRBs and pulsars using a unified methodological approach.
+Section 2 describes the source selection and data processing, while Section 3 and 4 presents and discusses the results.
 We conclude in Section 5.
-2 Sources selection and data processing
-2.1 Sources selection
-Figure 1: Isotropic-equivalent spectral luminosity versus the product of observing frequency and pulse width for various pulsars and FRBs.
-The brightest strong pulses from RRATs with relatively high $B_{\rm s}$ (J1846-0257 and 1854+0306) are represented by red and black filled stars, while their weak emissions are denoted by unfilled stars (Zhang et al., 2024).
-The brightest giant pulses from J0534+2200 (Crab pulsar) (Bera & Chengalur, 2019) and J0540--6919  (Geyer et al., 2021) are shown as blue and violet filled triangles, with their normal emissions  (Manchester et al., 2005) marked by unfilled triangles.
-The brightest burst from the nearby FRB 20200120E (Zhang et al., 2024a) is indicated by an orange filled square, with its relatively weaker bursts (Nimmo et al., 2023) are marked by unfilled squares.
-The brightest bursts from cosmological FRB 20201124A (Kirsten et al., 2024) and FRB 20220912A (Ould-Boukattine et al., 2024) are represented by black and gray filled circles, with their relatively weaker bursts (Xu et al., 2022; Zhang et al., 2023) shown as unfilled circles.
-The pulsar (Manchester et al., 2005) population is indicated in thistle.
-Grey dashed lines represent constant isotropic-equivalent energy release, while grey dotted lines indicate constant brightness temperature.
-The selected sources, presented in Figure 1, show the brightest strong pulses that are at least 2--3 orders of magnitude more energetic than their relatively weak emissions.
-The energies of the detected radio pulses from these sources span roughly 15 orders of magnitude.
-Specifically, the studied sources include: (1) FRB 20201124A and FRB 20220912A: These are the only two FRBs with up to 2000 hours of follow-up observations, during which ultra-luminous bursts with isotropic energy up to 1041 erg were detected (Kirsten et al., 2024; Ould-Boukattine et al., 2024), placing them among the most energetic FRBs observed.
-(2) FRB 20200120E: A nearby FRB that exhibited a burst storm (Nimmo et al., 2023) and an abnormal burst more than 42 times brighter than previously detected bursts at a similar frequency (Zhang et al., 2024a).
-(3) J0534+2200 (the Crab pulsar) and J0540--6919: These pulsars are known for producing giant pulses with the largest excesses relative to their normal emissions (Bera & Chengalur, 2019; Geyer et al., 2021; Wang et al., 2019).
-(4) RRATs J1846--0257 and J1854+0306: These RRATs have relatively high surface magnetic fields ($\sim 2\times 10^{13}$ G) and show weak emissions that are up to 687 and 512 times fainter than their strong individual pulses.
-We also analysed data from RRAT J1913+1330 (Zhang et al., 2024b), which exhibits extreme pulse-to-pulse modulation and an energy distribution spanning 3 orders of magnitude but could be well-fitted by two log-normal functions.
-Its distribution tail at the high energy part provides a useful comparison with the four source kinds above.
+2 Source selection and data processing
+2.1 Source selection
+Our selection prioritizes sources that exhibit flat high-energy tails in their burst energy distributions.
+Among FRBs, we include four cosmological sources with demonstrated power-law indices of $\alpha\gtrsim-1$ at high energies: FRB 20121102A, FRB 20201124A, FRB 20220912A, and FRB 20240114A.
+We also include the nearby FRB 20200120E, which shows a tentative high-energy tail ($\alpha\sim-0.98$), a key target of our study.
+For pulsars, we focus on J1846-0257 and J1854+0306, two RRATs exhibiting flat tails, in contrast to the steep energy distributions typically observed in giant pulse pulsars.
+For comparison, we selected the well-studied giant pulse pulsars J0534+2200 (the Crab pulsar) and J0540-6919, known for producing giant pulses with the largest excesses relative to their normal emissions.
+However, detailed burst properties for FRB 20220912A and FRB 20240114A are currently unavailable, limiting our ability to conduct comprehensive energy distribution analyses.
+Similarly, FRB 20200120E has only one documented ultra-bright burst, which restricts statistical analysis.
+Therefore, our quantitative energy distribution analysis focuses on the data of FRB 20121102A (Li et al., 2021, ; the largest available sample) and FRB 20201124A (Kirsten et al., 2024, ; the sample containing the brightest burst) among the FRB sources, J1846-0257 and J1854+0306 (Zhang et al., 2024a) among the pulsar sources, and J0534+2200 (Bera & Chengalur, 2019, the sample containing the largest number of bright bursts) and J0540-6919 (Geyer et al., 2021, ; the most extensive sample) as giant pulse comparison sources.
 2.2 Data processing
@@
 $\centering E_{S}=\frac{4\pi D_{L}^{2}}{1+z}\,F,\@add@centering$ (1)
-where $D_{L}$ is the luminosity distance, $z$ is the redshift, and $F$ is the pulse fluence. The energy values for each source were then normalized to their brightest bursts.
-Since only one bright burst has been detected from FRB 20200120E (Zhang et al., 2024a) and detailed burst properties for FRB 20220912A are currently unavailable (Ould-Boukattine et al., 2024), we were unable to reprocess their high-energy distributions.
-To compare our results with their reported distributions, which are cumulative, our initial analysis used a cumulative distribution, adjusting the bins for each source to cover a similar range in log-10 space.
-We also provide a differential distribution analysis.
+where $D_{L}$ is the luminosity distance, $z$ is the redshift, and $F$ is the reported pulse fluence from the burst samples described above.
+The energy values for each source were then normalized to their brightest bursts, as our analysis focuses on comparing the high-energy tails.
+Since only the cumulative energy distribution tail is available for the only nearby repeating FRB 20200120E, we used the power-law indices ($\alpha$) derived from the cumulative energy distribution in this study, which is given by:
+$\centering{\rm log}_{10}N(\geq E)=C+\alpha\,{\rm log}_{10}E,\@add@centering$ (2)
+where $N(\geq E)$ represents the predicted number of bursts with specific energy above $E$.
+In our distribution analysis, we adjusted the bins for each source to cover a similar range in log-10 space.
 To enable comparisons between different sources, we defined a density that is calculated as the number of bursts per bin divided by the total number of detected bursts.
-The power-law-like tails were fitted using the Python package curve_fit function from the SciPy library (Virtanen et al., 2020).
-For data with complex uncertainties, we performed the fitting within a Bayesian framework using Markov Chain Monte Carlo (MCMC) sampling, via the emcee Python package (Foreman-Mackey et al., 2013).
+For data of pulses from a single source (i.e., Figure 1), we considered only simple symmetric uncertainties in the counts, and fitted the power-law-like tails using the Python package curve_fit function from the SciPy library (Virtanen et al., 2020).
+The cutoff point for fitting was determined by identifying the location where the standard error was minimized.
+For data of pulses from different sources (i.e., Figure 2), we performed the fitting within a Bayesian framework using Markov Chain Monte Carlo (MCMC) sampling, via the emcee Python package (Foreman-Mackey et al., 2013).
+The fitting was conducted in log-log space using an asymmetric Gaussian likelihood.
 3 Results
@@
 Refer to caption Refer to caption Refer to caption
-Figure 2: Cumulative specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919.
+Figure 1: Cumulative specific energy distributions of high-energy bursts from FRB 20201124A and FRB 20121102A, single pulses from RRATs J1846-0257 and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540-6919.
 The energy values for each source are normalized to their respective brightest bursts.
 Each high-energy tail is fitted with a simple power law model.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 3: Differential specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919.
-The energy values for each source are normalized to their respective brightest bursts.
-Each high-energy tail is fitted with a simple power law model.
-Figure 4: Comparison of cumulative (top) and differential (bottom) normalized specific energy distributions of high-energy bursts from FRB 20201124A, single pulses from RRATs J1913+1330, J1846--0257, and J1854+0306, and giant pulses from the Crab pulsar (J0534+2200) and J0540--6919.
-The energy values for each source are normalized to their respective brightest bursts.
-The density represents the number of bursts per bin divided by the total number of detected bursts.
+The fitting cutoff points are indicated by black dashed lines.
 Figure 1 presents the cumulative specific energy distributions of high-energy bursts from the selected sources with relatively sufficient burst counts, most of which can be well-fitted by a simple power-law model.
-The high-energy tail of the distribution for RRAT J1913+1330 is forcefully fitted using a power-law function, although the fit is not as good as for the other sources, in order to enable comparison.
-Specifically, the fitted power-law indices ($\alpha$) for FRB 20201124A (46 bursts  [^2] For events detected simultaneously by multiple telescopes, we selected bursts from the observation with the largest bandwidth.), RRATs J1846--0257 (72 bright pulses) and J1854+0306 (325 bright pulses) are $-0.75\pm 0.14$, $-0.84\pm 0.10$ and $-0.76\pm 0.03$, respectively, all of which are $\gtrsim-1$.
-In contrast, the indices for the Crab pulsar ($-1.98\pm 0.07$, 1955 giant pulses), J0540--6919 ($-2.29\pm 0.07$, 865 bright single pulses) and RRAT J1913+1330 ($-3.04\pm 0.29$, 1955 individual pulses) are significantly steeper, at $\lesssim-2$.
-Although the available data for bursts from FRB 20200120E and FRB 20220912A are insufficient to reprocess their high-energy distributions, the reported fitted power-law indices are $-0.98\pm 0.06$ (Zhang et al., 2024a) for FRB 20200120E and range from $-0.99$ to $-0.60$ (Ould-Boukattine et al., 2024) for FRB 20220912A at around 1.4 GHz, both indicate relatively flat indices.
-Furthermore, we also present the differential specific energy distributions of high-energy bursts from these source in Figure 3.
-Similar to the cumulative distributions, the fitted indices using a power-law function for the differential distributions at the high energy tails ($\beta$) of FRB 20201124A ($-0.40\pm 0.21$), RRATs J1846--0257 ($-0.63\pm 0.16$) and J1854+0306 ($-0.65\pm 0.05$) are significantly flatter than those of the Crab pulsar ($-1.98\pm 0.09$), J0540--6919 ($-2.0\pm 0.1$) and RRAT J1913+1330 ($-2.72\pm 0.31$).
-However, the sample sizes for some sources, especially RRAT J1846--0257 and FRB 20201124A, are limited, and the fitted uncertainties are relatively large.
-As shown in the comparison in Figure 4, the energy distributions of FRB 20201124A, RRATs J1846--0257 and J1854+0306 decline much more slowly at high energies compared to those of the giant-pulse pulsars and RRAT J1913+1330.
+Specifically, the fitted power-law indices ($\alpha$) for FRB 20201124A (46 bursts [^3] For events detected simultaneously by multiple telescopes, we selected bursts from the observation with the largest bandwidth.), RRATs J1846-0257 (72 bright pulses) and J1854+0306 (325 bright pulses) are $0.67\pm 0.05$, $0.82\pm 0.03$ and $0.77\pm 0.01$, respectively, all of which are $\gtrsim-1$.
+In contrast, the indices for the Crab pulsar ($1.98\pm 0.07$, 1955 giant pulses) and J0540-6919 ($2.29\pm 0.1$, 865 bright single pulses) are significantly steeper, at $\lesssim-2$.
+The index for FRB 20121102 is $1.51\pm 0.04$ (1652 bursts), falling between the above two categories.
+Although the available data for bursts from FRB 20200120E and FRB 20220912A are insufficient to reprocess their high-energy distributions, the reported fitted power-law indices from cumulative distributions are $0.98\pm 0.06$ (Zhang et al., 2024b) for FRB 20200120E and range from 0.99-0.99 to 0.60-0.60 (Ould-Boukattine et al., 2024) for FRB 20220912A at around 1.4 GHz.
+Both values suggest relatively flat indices, consistent with the first group.
 4 Discussion
 4.1 Bright bursts of various radio sources
-Figure 5: Specific energy of the brightest bursts versus their detection rates for FRB 20201124A, FRB 20220912A, FRB 20200120E and the RRATs J1846--0257 and J1854+0306.
+Figure 2: Specific energy of the brightest bursts versus their detection rates for FRB 20201124A, FRB 20220912A, FRB 20200120E and the RRATs J1846-0257 and J1854+0306.
 The data are fitted with a power law relation (red dashed line) using a Bayesian framework, and the shaded region represents the 68% credible interval for the fit.
 Posterior distributions and a parameter correlation plot for the power law fit parameters are shown in the inset.
+Three horizontal contour lines are added at 100 hours, 1000 hours, and 10000 hours observation time, with their intersections with each source’s distribution line indicating the expected rate and energy of the brightest bursts detectable for these observation durations.
 Sources with flat energy distributions ($\alpha\gtrsim-1$) at high energies suggest that their “bright pulses” might contribute significantly to the total radio pulse energy release.
 Therefore, the detection of the brightest bursts from these sources could provide insight into a potential connection between cosmological high-luminosity radio bursts and nearby low-luminosity ones.
-Figure 5 presents the specific energy of the brightest bursts and the corresponding detection rates for the studied sources with flat energy distributions, i.e., RRATs J1846--0257 and J1854+0306, FRB 20200120E, FRB 20201124A and FRB 20220912A.
-Interestingly, these data points align well with a simple power-law fit of $\alpha=-0.26\pm 0.05$.
-Due to the sample size of only five points and the stochastic nature of detecting the “brightest burst”, we remain cautious about interpreting this correlation as a physical relationship.
-Nevertheless, this trend suggests that high-luminosity radio bursts, such as FRBs, could be naturally derived from low-luminosity radio bursts, such as those from pulsars.
-Although an extremely shallow power-law index is implied, the flatter energy distributions at higher energies observed in these sources may provide clues for understanding this tentative connection.
+Figure 2 presents the specific energy of the brightest bursts and the corresponding detection rates [^4] Here, we estimate the rate based on a single detected event, with a $1\sigma$ confidence interval from 0.173 to 3.3 (Gehrels, 1986).
+for the studied sources with flat energy distributions, i.e., RRATs J1846-0257 and J1854+0306, FRB 20200120E, FRB 20201124A and FRB 20220912A.
+Interestingly, these data points could align with a power-law fit of $\alpha_{\rm Bri}=-0.26\pm 0.05$.
+Due to the sample size of only five points and the stochastic nature of detecting the “brightest burst”, which usually depends on the telescope sensitivities and monitoring lengths, we will not suggest interpreting this correlation as a physical relationship at this stage.
+Three horizontal contours at 100, 1000, and 10000 hours mark the expected rate and energy of the brightest detectable bursts at each observation duration.
+Even extending to 10,000 h, the expected brightest RRAT energies lie below nearby FRBs, and nearby FRBs lie below bright cosmological FRBs.
+This trend suggests that a much flatter index than that observed in current sources such as FRB 20200120E is required to naturally derive high-luminosity radio bursts, such as FRBs, from low-luminosity bursts, such as those from pulsars.
+It is worth noting that the power-law indices used here were derived from cumulative distributions, in which the data points are not independent.
+As a result, our line fitting may produce biased results and underestimate the associated uncertainties.
+This limitation might be further addressed through more robust methods, such as the maximum likelihood approach introduced by James et al.
+(2019), or through future detections of a sufficient number of bright bursts to allow comparison using differential distributions.
 4.2 Survey strategies for nearby FRBs
-As presented in Figure 2 and 3 of Section 3, the fitting of the cumulative energy distribution at high energies using a power-law function yields better results with smaller uncertainties compared to the differential distributions.
-This difference may arise because the limited burst numbers for the studied sources are insufficient to provide a reliable sample for the differential distributions.
-Therefore, we used the power-law indices ($\alpha$) derived from the cumulative energy distribution to estimate the detection rate in this section, which is given by:
-$\centering{\rm log}_{10}N(\geq E)=C+\alpha\,{\rm log}_{10}E,\@add@centering$ (2)
-where $N(\geq E)$ represents the predicted number of bursts with specific energy above $E$.
-For a telescope operating at a fluence completeness threshold of $F_{\rm det}$, bursts with energy above the threshold $E_{\rm det}=4\pi D_{\rm det}^{2}F_{\rm det}$ are expected to be detected, where $D_{\rm det}$ is the distance to the target source, and assuming a detected burst rate of $R_{\rm det}$.
+For a telescope operating at a fluence completeness threshold of $F_{\rm det}$, bursts with energy above the threshold $E_{\rm det}=4\pi D_{\rm det}^{2}F_{\rm det}$ are expected to be detected, where $D_{\rm det}$ is the distance to the target source, and assuming a reference burst rate of $R_{\rm ref}$.
 Applying Equation 2, the burst rate expected for bursts detectable by an arbitrary radio telescope with a fluence threshold $F_{\rm th}$, for a source at a distance of $D_{\rm tar}$, can be written as:
-$\centering R_{\rm th}\approx R_{\rm det}\times\left(\frac{F_{\rm th}}{F_{\rm det}}\right)^{\alpha}\left(\frac{D_{\rm tar}}{D_{\rm det}}\right)^{2\alpha}.\@add@centering$ (3)
+$\centering R_{\rm th}\approx R_{\rm ref}\times\left(\frac{F_{\rm th}}{F_{\rm det}}\right)^{\alpha}\left(\frac{D_{\rm tar}}{D_{\rm det}}\right)^{2\alpha}.\@add@centering$ (3)
+As discussed by Zhang et al.
+(2024b), current FRB surveys may suffer from selection effects, potentially missing FRBs with relatively low luminosity and narrow widths from nearby GC systems.
+For example, the current CHIME system only records baseband data for real-time detections with high signal-to-noise ratios (typically between 10 and 12), and at a time resolution of approximately 1 ms (CHIME/FRB Collaboration et al., 2024).
+Such a system would only be sensitive to a burst similar to the brightest one from FRB 20200120E (with an effective width of $\leq 0.161$ ms and specific luminosity of $\sim 3.0\times 10^{30}$ erg s-1 Hz-1) at distances around $\sim$10 Mpc, and the event rate is extremely low (Zhang et al., 2024b).
+For more frequent, lower-luminosity bursts from FRB 20200120E, the system is realistically sensitive only to its current distance of $\sim 3$ Mpc.
 Kremer et al.
-(2023) presented prospects for detecting FRBs from GC in nearby galaxies using FAST (Jiang et al., 2020) and MeerKAT (Bailes et al., 2020), suggesting that NGC 4486 (M87) could be the best candidate target.
+(2023) presented prospects for detecting FRBs from GCs in nearby galaxies using FAST and MeerKAT, suggesting that NGC 4486 (M87) could be the best candidate target.
 However, the fluence threshold ($F_{\rm th}$) of a telescope may be significantly affected by the continuum radio flux densities ($S_{\rm tar}$) from the target galaxies (Suresh et al., 2021), which was not thoroughly discussed in detection rate estimation by Kremer et al.
@@
 Therefore, for telescopes with high sensitivities, galaxies with relatively small $S_{\rm tar}$, such as NGC 4649 (M60), which has $S_{\rm tar}\sim 0.03$ at 1.4 GHz (Brown et al., 2011) and is estimated to contain approximately 4000 GCs (Harris et al., 2013), could be more efficient search targets.
-At a given distance, the number of potential sources is proportional to the field of view (FOV) of the telescope. Combining the FOV with Equation 3, the detection rate of the telescope is given by:
+At a given distance, assuming an isotropic source distribution, the number of potential sources is proportional to the field of view (FOV) of the telescope.
+Combining the FOV with Equation 3, the detection rate of the telescope is given by:
 $\centering R\propto F_{\rm th}^{\alpha}\cdot D_{\rm tar}^{2\alpha}\cdot{\rm FOV}.\@add@centering$ (5)
@@
 Therefore, in addition to telescopes with high sensitivity, such as FAST and MeerKAT, observations with large FOV but relatively lower sensitivity may also be efficient for capturing such events.
-Figure 6: Parkes cryoPAF 63 pointings covering more than 10,000 valid GCs.
-The two most suggested pointings, TP0 (Dec $\sim$ 13 deg) and TP1 (Dec $\sim$ 8 deg), are marked by black and red stars, respectively, with their corresponding sky coverage shown by circles.
+Figure 3: Parkes cryoPAF 1,832 pointings covering more than 7,000 clean GCs, clustered into six groups.
+The sky coverage of the most recommended pointing in each group, TP0 (RA = 187.2 deg; Dec = 13.2 deg), TP1 (RA = 186.6 deg; Dec = 8.1 deg), TP2 (RA = 225.9 deg; Dec = 1.5 deg), TP3 (RA = 55.0 deg; Dec = -18.6 deg), TP4 (RA = 54.6 deg; Dec = -35.1 deg) and TP5 (RA = 157.4 deg; Dec = -35.5 deg), is indicated by black circle.
 Table 1: Properties of the 24 GC systems covered by Parkes cryoPAF pointing TP0 (RA = 187.2 deg; Dec = 13.2 deg).
@@
 (2010).
-Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm V\_GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
+Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm C,GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
 NGC4486 (M87) 187.706 12.391 17.0 0.95 13000 6500 $\sim$ 210 (a)
 NGC4374 (M84) 186.265 12.887 18.5 0.96 4301 4301 $\sim$ 6.1 (b)
-NGC4406 (M86) 186.549 12.946 17.1 0.68 2800 2800 $<$ 0.1 (c)
-NGC4473 187.454 13.429 15.2 0.34 376 376 $<$ 0.1 (a)
-NGC4435 186.919 13.079 16.6 0.30 345 345 $<$ 0.1 (c)
-NGC4459 187.250 13.979 16.0 0.78 218 218 $<$ 0.1 (d)
-NGC4474 187.473 14.069 15.5 0.91 116 116 $<$ 0.1 (c)
-NGC4458 187.240 13.242 16.3 0.06 72 72 $<$ 0.1 (e)
-NGC4387 186.424 12.810 18.0 0.85 70 70 $<$ 0.1 (e)
-VCC-940 186.696 12.454 18.7 0.89 61 61 $<$ 0.1 (c)
-NGC4479 187.577 13.578 17.4 0.53 59 59 $<$ 0.1 (c)
-NGC4478 187.574 12.327 17.0 0.95 58 58 $<$ 0.1 (c)
-VCC-1386 187.964 12.657 16.0 0.92 27 27 $<$ 0.1 (c)
-NGC4476 187.496 12.349 17.7 0.90 20 20 $<$ 0.1 (c)
-VCC-871 186.523 12.560 16.0 0.92 18 18 $<$ 0.1 (c)
-VCC-965 186.763 12.561 16.9 0.77 15 15 $<$ 0.1 (c)
-VCC-1185 187.348 12.451 16.9 0.76 14 14 $<$ 0.1 (c)
-NGC4431 186.864 12.290 15.8 0.97 11 11 $<$ 0.1 (c)
-VCC-1104 187.117 12.824 15.7 0.38 8 8 $<$ 0.1 (c)
-VCC-896 186.594 12.787 16.0 0.72 5 5 $<$ 0.1 (c)
-VCC-1105 187.114 14.156 16.0 0.96 5 5 $<$ 0.1 (c)
+NGC4406 (M86) 186.549 12.946 17.1 0.68 2800 2800 << 0.1 (c)
+NGC4473 187.454 13.429 15.2 0.34 376 376 << 0.1 (a)
+NGC4435 186.919 13.079 16.6 0.30 345 345 << 0.1 (c)
+NGC4459 187.250 13.979 16.0 0.78 218 218 << 0.1 (d)
+NGC4474 187.473 14.069 15.5 0.91 116 116 << 0.1 (c)
+NGC4458 187.240 13.242 16.3 0.06 72 72 << 0.1 (e)
+NGC4387 186.424 12.810 18.0 0.85 70 70 << 0.1 (e)
+VCC-940 186.696 12.454 18.7 0.89 61 61 << 0.1 (c)
+NGC4479 187.577 13.578 17.4 0.53 59 59 << 0.1 (c)
+NGC4478 187.574 12.327 17.0 0.95 58 58 << 0.1 (c)
+VCC-1386 187.964 12.657 16.0 0.92 27 27 << 0.1 (c)
+NGC4476 187.496 12.349 17.7 0.90 20 20 << 0.1 (c)
+VCC-871 186.523 12.560 16.0 0.92 18 18 << 0.1 (c)
+VCC-965 186.763 12.561 16.9 0.77 15 15 << 0.1 (c)
+VCC-1185 187.348 12.451 16.9 0.76 14 14 << 0.1 (c)
+NGC4431 186.864 12.290 15.8 0.97 11 11 << 0.1 (c)
+VCC-1104 187.117 12.824 15.7 0.38 8 8 << 0.1 (c)
+VCC-896 186.594 12.787 16.0 0.72 5 5 << 0.1 (c)
+VCC-1105 187.114 14.156 16.0 0.96 5 5 << 0.1 (c)
 NGC4486B 187.633 12.490 16.3 0.83 4 4 $\sim$ 0.4 (c)
-VCC-1272 187.567 13.308 16.0 0.37 3 3 $<$ 0.1 (c)
-VCC-1077 187.043 12.807 16.0 0.42 2 2 $<$ 0.1 (c)
-Table 2: Properties of the 12 GC systems covered by Parkes cryoPAF pointing TP1 (RA = 186.6 deg; Dec = 8.1 deg).
-$d_{\rm L}$ and $D_{\rm TP}$ represent the distance from the GC system centre to the earth and the pointing, respectively.
+VCC-1272 187.567 13.308 16.0 0.37 3 3 << 0.1 (c)
+VCC-1077 187.043 12.807 16.0 0.42 2 2 << 0.1 (c)
+Table 2: Properties of the GC systems covered by Parkes cryoPAF pointings: 12 systems by TP1 (RA = 186.6 deg; Dec = 8.1 deg), 3 systems by TP2 (RA = 225.9 deg; Dec = 1.5 deg), 2 systems by TP3 (RA = 55.0 deg; Dec = -18.6 deg), 21 systems by TP4 (RA = 54.6 deg; Dec = -35.1 deg), and 2 systems by TP5 (RA = 157.4 deg; Dec = -35.5 deg).
 References for the galaxy continuum radio flux densities at 1.4 GHz ($S_{\rm tar,1.4GHz}$) are: (a) White & Becker (1992), (b) Brown et al.
-(2011), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/).
-Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm V\_GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
-NGC4472 (M49) 187.445 8.000 17.0 0.84 7000 7000 $\sim$ 0.8 (a)
-NGC4365 186.118 7.318 23.3 0.92 3246 3246 $<$ 0.1 (b)
-NGC4434 186.903 8.154 22.5 0.30 141 141 $<$ 0.1 (c)
-VCC-1254 187.521 8.074 16.0 0.91 26 26 $<$ 0.1 (c)
-VCC-1107 187.126 7.325 16.0 0.93 25 25 $<$ 0.1 (c)
-NGC4464 187.339 8.157 15.8 0.73 25 25 $<$ 0.1 (c)
-NGC4318 185.680 8.198 22.1 0.92 18 18 $<$ 0.1 (c)
-VCC-747 186.199 8.992 18.5 0.98 16 16 $<$ 0.1 (c)
-VCC-1167 187.301 7.879 16.0 0.73 14 14 $<$ 0.1 (c)
-VCC-571 185.671 7.950 23.8 0.93 11 11 $<$ 0.1 (c)
-VCC-1049 186.978 8.090 16.0 0.37 8 8 $<$ 0.1 (c)
-VCC-992 186.828 8.213 16.0 0.25 4 4 $<$ 0.1 (c)
+(2011), (c) estimated from National Radio Astronomy Observatory Very Large Array Sky Survey (NVSS, https://www.cv.nrao.edu/nvss/), (d) Condon et al.
+(1998) and (e) Allison et al.
+(2014).
+Pointing Name RA (deg) Dec (deg) $d_{\rm L}$ (Mpc) $D_{\rm TP}$ (deg) $N_{\rm GC}$ $N_{\rm C,GC}$ $S_{\rm tar,1.4GHz}$ (Jy)
+TP1 NGC4472 (M49) 187.445 8.000 17.0 0.84 7000 7000 $\sim$ 0.8 (a)
+NGC4365 186.118 7.318 23.3 0.92 3246 3246 << 0.1 (b)
+NGC4434 186.903 8.154 22.5 0.30 141 141 << 0.1 (c)
+VCC-1254 187.521 8.074 16.0 0.91 26 26 << 0.1 (c)
+VCC-1107 187.126 7.325 16.0 0.93 25 25 << 0.1 (c)
+NGC4464 187.339 8.157 15.8 0.73 25 25 << 0.1 (c)
+NGC4318 185.680 8.198 22.1 0.92 18 18 << 0.1 (c)
+VCC-747 186.199 8.992 18.5 0.98 16 16 << 0.1 (c)
+VCC-1167 187.301 7.879 16.0 0.73 14 14 << 0.1 (c)
+VCC-571 185.671 7.950 23.8 0.93 11 11 << 0.1 (c)
+VCC-1049 186.978 8.090 16.0 0.37 8 8 << 0.1 (c)
+VCC-992 186.828 8.213 16.0 0.25 4 4 << 0.1 (c)
+TP2 NGC5846 226.622 1.606 24.9 0.73 4670 4670 << 0.1 (b)
+NGC5813 225.297 1.702 32.2 0.64 2900 2900 << 0.1 (b)
+NGC5845 226.503 1.634 25.9 0.62 180 180 << 0.1 (c)
+TP3 NGC1407 55.049 -18.580 28.8 0.05 7000 7000 << 0.1 (d)
+NGC1400 54.879 -18.688 26.4 0.14 2960 2960 << 0.1 (b)
+TP4 NGC1399 54.621 -35.451 20.7 0.35 6000 6000 $\sim$ 0.6 (e)
+NGC1404 54.716 -35.594 20.4 0.50 770 770 << 0.1 (b)
+NGC1427 55.581 -35.393 20.5 0.85 500 500 << 0.1 (b)
+NGC1380 54.115 -34.976 18.9 0.42 424 424 << 0.1 (b)
+NGC1387 54.238 -35.507 19.8 0.50 390 390 << 0.1 (b)
+NGC1374 53.819 -35.226 19.6 0.65 360 360 << 0.1 (b)
+NGC1379 54.016 -35.441 20.0 0.59 225 225 << 0.1 (b)
+NGC1380B 54.287 -35.195 22.0 0.27 156 156 << 0.1 (b)
+NGC1375 53.820 -35.266 20.0 0.66 86 86 << 0.1 (c)
+NGC1381 54.132 -35.295 18.8 0.43 71 71 << 0.1 (c)
+NGC1380A 54.198 -34.740 16.9 0.49 70 70 << 0.1 (c)
+ESO-358-G2 53.879 -34.447 20.8 0.88 60 60 << 0.1 (c)
+FCC-182 54.226 -35.375 19.6 0.41 59 59 << 0.1 (c)
+NGC1389 54.299 -35.746 20.9 0.69 48 48 << 0.1 (c)
+NGC1428 55.596 -35.153 20.7 0.82 42 42 << 0.1 (c)
+NGC1427A 55.039 -35.625 16.0 0.63 38 38 << 0.1 (c)
+ESO-358-G4 54.538 -34.519 20.7 0.58 30 30 << 0.1 (c)
+FCC-136 53.623 -35.546 19.7 0.91 25 25 << 0.1 (c)
+FCC-254 55.253 -35.743 20.4 0.83 6 6 << 0.1 (c)
+FCC-189 54.284 -34.731 20.0 0.45 1 1 << 0.1 (c)
+FCC-146 53.798 -35.323 20.0 0.69 1 1 << 0.1 (c)
+TP5 NGC3258 157.223 -35.605 32.1 0.18 6000 6000 << 0.1 (b)
+NGC3268 157.503 -35.326 34.8 0.19 4750 4750 << 0.1 (b)
 A cryogenic phased array receiver (cryoPAF) is currently being completed and tested on the Parkes telescope.
-The new cryoPAF is expected to have a system temperature of $<$ 20 K and a gain of $0.735{\rm K/Jy}$, resulting in a corresponding SEFD of $S_{\rm Parkes}=27.21$ Jy.
+The new cryoPAF is expected to have a system temperature of << 20 K and a gain of $0.735{\rm K/Jy}$, resulting in a corresponding SEFD of $S_{\rm Parkes}=27.21$ Jy.
 Its wide FOV spans up to 3 square degrees and consists of 72 adjacent beams, covering a frequency range of 700-1950 MHz  [^6] As cryoPAF is still being tested, the accurate parameters of the full-beam system performance are unavailable for now.
 We used the values referenced from the Parkes cryoPAF introduction by Alex Dunning: https://research.csiro.au/ratechnologies/wp-content/uploads/sites/295/2022/11/PAFAR2022-Dunning-CryoPAF_for_Parkes.pdf.
-Based on the cryoPAF’s sky coverage, which is modelled as a circle with a radius of approximately 1 deg in this work, and the catalogue of GC systems (Harris et al., 2013) within a relatively close distance ($<$ 40 Mpc), we conducted a grid analysis of the sky with a 0.1 deg step size to estimate the number of GCs covered by the cryoPAF FOV.
+Based on the cryoPAF’s sky coverage, which is modelled as a circle with a radius of approximately 1 deg in this work, and the catalogue of GC systems (Harris et al., 2013) within a relatively close distance (<< 40 Mpc), we conducted a grid analysis of the sky with a 0.1 deg step size to estimate the number of GCs covered by the cryoPAF FOV.
 Notably, about 20% of GCs lie within a 4-arcminute radius of the centre of M87 (Strader et al., 2011), where the radio background is comparable to $S_{\rm Parkes}$ (Perley & Butler, 2017).
-Given this, the cryoPAF observation should cover the majority of GCs in M87 with adequate sensitivity, and we adopt a conservative estimate that 50% of the GCs are valid targets for Parkes observation (denoted as $N_{\rm V\_GC}$).
-As shown in Figure 6, among the more than 6,000,000 analysed pointings, 63 could cover more than 10,000 valid GCs, all with Dec near 13 or 8 deg.
-Pointings TP0 (RA = 187.2 deg; Dec = 13.2 deg) and TP1 (RA = 186.6 deg; Dec = 8.1 deg) cover the largest number of valid GCs in these two Dec ranges.
-Based on this, we suggest these two pointings as the best targets for searching for FRBs in nearby GCs.
-As listed in Tables 1 and 2, the Parkes cryoPAF pointing TP0 could cover 24 GC systems, expected to include 21,607 GCs, of which 15,107 are valid targets.
-TP1 could cover 12 GC systems, expected to include 10,535 GCs, all of which are valid.
+Given this, the cryoPAF observation should cover the majority of GCs in M87 with adequate sensitivity, and we adopt a conservative estimate that 50% of the GCs are clean targets for Parkes observation (denoted as $N_{\rm C,GC}$).
+As shown in Figure 3, among the more than 6,000,000 analysed pointings, 1,832 could cover more than 7,000  [^7] We set this threshold based on the estimated number of $N_{\rm C,GC}$ in GC systems within 40 Mpc.
+clean GCs, clustered into six groups.
+Based on both the number of $N_{\rm C,GC}$ within the sky coverage and the distance from the pointing to GC system centres, the most recommended pointings for each group are: TP0 (RA = 187.2 deg; Dec = 13.2 deg), TP1 (RA = 186.6 deg; Dec = 8.1 deg), TP2 (RA = 225.9 deg; Dec = 1.5 deg), TP3 (RA = 55.0 deg; Dec = -18.6 deg), TP4 (RA = 54.6 deg; Dec = -35.1 deg) and TP5 (RA = 157.4 deg; Dec = -35.5 deg).
+Full details of all 1,832 pointings are provided in the Supplementary Materials.
+As listed in Tables 1 and 2, Parkes cryoPAF pointing TP0 could cover 24 GC systems, expected to include 21,607 GCs, of which 15,107 are clean targets.
+TP1 covers 12 GC systems with 10,535 $N_{\rm C,GC}$; TP2 covers 3 GC systems with 7,750 $N_{\rm C,GC}$; TP3 covers 2 GC systems with 9,960 $N_{\rm C,GC}$; TP4 covers 21 GC systems with 9,361 $N_{\rm C,GC}$; and TP5 covers 2 GC systems with 10,750 $N_{\rm C,GC}$.
 Table 3: Predicted FRB detection rates in nearby GC systems by FAST, MeerKAT and Parkes cryoPAF.
 FRB detection rate (hr-1)
-Target Instrument $N_{\rm GC}$ $N_{\rm V\_GC}$ $\alpha$ = --0.2 $\alpha$ = --0.5 $\alpha$ = --1.0 $\alpha$ = --1.5 $\alpha$ = --2.0
-M60 FAST 4000 1203 0.16(0.01--0.48) 0.35(0.02--1.06) 1.36(0.07--4.08) 5.21(0.27--15.64) 19.98(1.02--59.93)
-M60 MeerKAT 4000 4000 0.37(0.02--1.11) 0.49(0.02--1.46) 0.77(0.04--2.31) 1.21(0.06--3.64) 1.92(0.1--5.75)
-TP0 Parkes cryoPAF 21607 15107 1.3(0.07--3.89) 1.53(0.08--4.6) 2.03(0.1 --6.1) 2.72(0.14--8.15) 3.65(0.19--10.96)
-TP1 Parkes cryoPAF 10535 10535 0.88(0.04--2.65) 1.01(0.05--3.03) 1.28(0.07--3.84) 1.65(0.08--4.94) 2.15(0.11--6.44)
-[…]
+Target Instrument $N_{\rm GC}$ $N_{\rm C,GC}$ $\alpha$ = -0.2 $\alpha$ = -0.5 $\alpha$ = -1.0 $\alpha$ = -1.5 $\alpha$ = -2.0
+M60 FAST 4000 1203 0.16(0.01-0.48) 0.35(0.02-1.06) 1.36(0.07-4.08) 5.21(0.27-15.64) 19.98(1.02-59.93)
+M60 MeerKAT 4000 4000 0.37(0.02-1.11) 0.49(0.02-1.46) 0.77(0.04-2.31) 1.21(0.06-3.64) 1.92(0.1-5.75)
+TP0 Parkes cryoPAF 21607 15107 1.3(0.07-3.89) 1.53(0.08-4.6) 2.03(0.1 -6.1) 2.72(0.14-8.15) 3.65(0.19-10.96)
+TP1 Parkes cryoPAF 10535 10535 0.88(0.04-2.65) 1.01(0.05-3.03) 1.28(0.07-3.84) 1.65(0.08-4.94) 2.15(0.11-6.44)
+TP2 Parkes cryoPAF 7750 7750 0.56(0.03-1.68) 0.51(0.03-1.54) 0.45(0.02-1.34) 0.40(0.02-1.19) 0.35(0.02-1.06)
+TP3 Parkes cryoPAF 9960 9960 0.71(0.04-2.14) 0.64(0.03-1.92) 0.54(0.03-1.61) 0.45(0.02-1.35) 0.38(0.02-1.13)
+TP4 Parkes cryoPAF 9361 9361 0.76(0.04-2.28) 0.82(0.04-2.47) 0.94(0.05-2.82) 1.07(0.05-3.22) 1.23(0.06-3.69)
+TP5 Parkes cryoPAF 10750 10750 0.72(0.04-2.16) 0.58(0.03-1.75) 0.41(0.02-1.24) 0.29(0.01-0.87) 0.21(0.01-0.62)
+Furthermore, we estimated the predicted FRB detection rates for FAST and MeerKAT targeting M60, as well as for Parkes cryoPAF targeting TP0 and TP1, using power-law indices $\alpha=-0.2,\,-0.5,\,-1.0,\,-1.5$ and 2.0-2.0.
+Most of the essential values used in this calculation are similar to Kremer et al.
+(2023).
+The fluence thresholds of FAST and MeerKAT are 0.015 Jy ms (Niu et al., 2021) and 0.09 Jy ms (Bailes et al., 2020).
+The reference values for the reference rate ($R_{\rm ref}\sim 0.07$ hr-1), fluence threshold ($F_{\rm det}\sim 5$ Jy ms), and source distance ( $D_{\rm det}\sim 3.63$ Mpc) were derived from CHIME’s observations of FRB 20200120E (CHIME/FRB Collaboration et al., 2021; Bhardwaj et al., 2021; Kirsten et al., 2022).
+The detection of a single FRB repeater by CHIME in the GCs within the M81 region suggests a specific abundance of $N_{\rm obs}\sim 1/900\approx 1.1\times 10^{-3}$ (Bhardwaj et al., 2021; Harris et al., 2013; Kremer et al., 2023), with a 90% confidence range of $(5.6\times 10^{-3},\,3.3\times 10^{-3})$, incorporating the Poisson probability (Lu et al., 2022).
+For GC systems with $S_{\rm tar,1.4GHz}<0.1$ Jy, which are negligible compared to the SEFDs of all three telescopes, we ignored their effect on the minimum detectable fluence.
+For GCs with larger $S_{\rm tar,1.4GHz}$, we used the effective minimum detectable fluence calculated using Equation 4.
 As presented in Table 3, for steep power-law indices ($\alpha=-1.5,\,-2.0$), FAST is the most powerful telescope for detecting nearby FRBs, with a detection rate of up to 20 hr-1.
@@
 5 Conclusion
-Before the discovery of a population of FRBs (Thornton et al., 2013) that supported the astrophysical origin of the “Lorimer burst” (Lorimer et al., 2007), pulsars (Manchester et al., 2005) were considered the primary sources of bright, short-duration ($\lesssim$ second) radio pulses, which require coherent emission (Pietka et al., 2015).
-While FRBs also necessitate coherent emission, their luminosities are more than ten orders of magnitude higher.
-Initially, the apparent one-off nature of FRBs led to favouring catastrophic scenarios (Zhang, 2023), distinguishing them from pulsars.
-However, the discovery of repeating FRBs (Spitler et al., 2016), and the investigation that some Galactic radio pulsing sources might be one-off (Keane, 2016; Zhang & Yang, 2024), have reduced this distinction.
-Further observations of fainter FRB bursts (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Nimmo et al., 2023; Zhang et al., 2024a) and brighter pulses from magnetars or pulsars (Kirsten et al., 2021; Good & Chime/Frb Collaboration, 2020; Bera & Chengalur, 2019) suggest no clear gap between the isotropic-equivalent energy releases of various radio pulses (Zhang et al., 2024a).
-Additionally, polarimetric similarities between FRBs and pulsars have been observed (Feng et al., 2022; Niu et al., 2024).
-These findings highlight the importance of exploring potential connections between FRBs and pulsars, at least for certain sources.
-In this study, we have presented similarities in the high-energy distributions of two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB 20200120E, and two pulsars (RRATs J1846--0257 and J1854+0306).
+In this study, we have presented similarities in the high-energy distributions of two cosmological repeating FRBs (FRB 20201124A and FRB 20220912A), one nearby FRB 20200120E, and two pulsars (RRATs J1846-0257 and J1854+0306).
 All these sources exhibit power-law indices of $\alpha\gtrsim-1$, suggesting that their “bright pulses” might significantly contribute to the total radio pulse energy release.
-Furthermore, the brightest bursts from these sources fit well with a simple power-law model of $\alpha=-0.26\pm 0.05$.
-However, considering the limited sample size, we advise caution in interpreting this correlation as physical.
+Furthermore, the brightest bursts from these sources could be fitted with a power-law model of $\alpha_{\rm Bri}=-0.26\pm 0.05$.
+Although we will not suggest interpreting this correlation as a physical relationship, the trend suggests that an extremely flat index is required to naturally derive high-luminosity radio bursts, such as FRBs, from low-luminosity bursts, such as those from pulsars.
 We encourage further observations to expand the sample of high-energy bursts from both FRB-like and pulsar-like sources.
-Current FRB surveys may be limited by selection effects, potentially missing FRBs with relative low-luminosity from the nearby GC systems (Zhang et al., 2024a).
-Survey strategies will need to vary for sources with different energy distributions.
-In this work, we have discussed detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF for FRB in nearby GCs and suggested optimal survey targets.
-In principle, conducting combining observations for these targets using FAST for about $3$ hours and Parkes cryoPAF for $10-20$ hours could potentially lead to the discovery of new FRBs in nearby GCs, in practice.
-Additionally, a lack of detection after even a few hours of FAST observation can still provide strict constraints on flat $\alpha$ values.
+As discussed in Section 4.2, current FRB surveys may miss FRBs with relatively low luminosity and narrow widths from nearby GC systems.
+Survey strategies will need to vary to search for sources with different energy distributions.
+In this work, we have discussed detailed survey strategies for FAST, MeerKAT and Parkes cryoPAF for FRB searches in nearby GCs and suggested optimal survey targets.
+In principle, conducting combined observations for these targets using FAST for about $3$ hours and Parkes cryoPAF for 102010-20 hours could potentially lead to the discovery of new FRBs in nearby GCs, in practice.
+Additionally, a lack of detection after even a few hours of FAST observation can still provide strict constraints on the combination of $\alpha$ values and intrinsic activity rates, or FRB rates.
 Acknowledgments
-This research has been partially funded by the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos.
+We thank the reviewer for the valuable and detailed comments, which have greatly improved the quality of this manuscript.
+This research has been partially funded by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the National SKA Program of China (2022SKA0130100), the National Natural Science Foundation of China (grant Nos.
 12041306,12273113,12233002,12003028,12321003), the CAS Project for Young Scientists in Basic Research (Grant No.
-YSBR-063), the International Partnership Program of Chinese Academy of Sciences for Grand Challenges (114332KYSB20210018), the ACAMAR Postdoctoral Fellow, China Postdoctoral Science Foundation (grant No.
-2020M681758), and the Natural Science Foundation of Jiangsu Province (grant Nos.
-BK20210998).
+YSBR-063), , and the ACAMAR Postdoctoral Fellow.
+Supplementary Materials
+Full details of all 1,832 Parkes cryoPAF pointings, which cover more than 7,000 clean GCs, are provided in the file “Parkes_PAF_pointing_above7000.txt”.